OpenStudy (anonymous):

Solve: 5^(2x-3)=7^(x+1)

OpenStudy (callisto):

Have you learnt logarithm yet?

OpenStudy (anonymous):

yes. can you just check my answer?

OpenStudy (callisto):

Okay.

OpenStudy (anonymous):

log875/(log25-log7)

OpenStudy (callisto):

I got the same answer as yours.

OpenStudy (anonymous):

okay thanks. actually can you show me how you did that. I took long time to figure this out and I want to know if there is an easier way

OpenStudy (callisto):

I guess more or less the same as you did.. \[5^{2x-3}=7^{x+1}\]\[(2x-3)\log5=(x+1)\log7\]\[2x\log5-3\log5=x\log7+\log7\]\[x(2\log5-\log7)=\log7+3\log5\]\[x=\frac{\log7+3\log5}{(2\log5-\log7)}=\frac{\log(7\times5^3)}{\log25-\log7}\]

OpenStudy (anonymous):

I did not it that way but I guess that works too

OpenStudy (callisto):

How did you do it?

OpenStudy (anonymous):

\[5 ^{2x-3}=7^{x+1}\]\[\frac{ 5^{2x} }{ 5^{3} }=7^{x}*7\]\[5^{x}=7^{x}*875\]\[(\frac{ 25 }{ 7 })^{x}=875\]\[x=\log_{25/7}875=\frac{ \log875 }{ \log25-\log7 } \]

OpenStudy (anonymous):

it's so difficult to think of ways to solve math problems. The method always seems to escape me

OpenStudy (anonymous):

thanks for your help

OpenStudy (callisto):

Your method looks nice :) I was taught to solve question in that way :(